# Learning Whole-Body Loco-Manipulation for Omni-Directional Task Space Pose Tracking with a Wheeled-Quadrupedal-Manipulator

### v1

@@ IV-A RP: Reward Prioritization
where $d_{\theta}=||\log(R_{ee}^{*}R_{ee}^{T})||_{F}$ is the Frobenius norm of the logarithm of the special orthogonal group SO(3) and $d_{p}=||P_{ee}^{*}-P_{ee}||_{2}$ is Euclidean distance [24].
@@ IV-C Loco-Mani Fusion: Achieving Smooth Loco-Manipulation
$\epsilon_{t}^{\text{ref}}=\min\left\{\epsilon_{0}-vt,0\right\}$ (11)
@@
Action rate $\left|a_{t-1}-a_{t}\right|^{2}$0.003
Action smoothness $\left|a_{t-2}-2a_{t-1}+a_{t}\right|^{2}$0.001
[…]
Weighted Power $\sum\|\tau\cdot\dot{q}\|\cdot\Omega_{p}$ $-3.3\times 10^{-4}$
Weighted Torque $\sum\tau^{2}\cdot\Omega_{\tau}$ $-4\times 10^{-5}$
Relative Torques rate $\sum\frac{\|\tau_{t}-\tau_{t-1}|}{\tau_{\text{limit}}}$ $-0.1$
@@
$e^{\text{ref}}_{t}$ $=\min\left\{|\epsilon_{t}^{\text{ref}}-\epsilon_{t}|-\gamma,0\right\}$ (13)
@@ V-B Simulation: Ablation Studies
Success Rate $\uparrow$ Aver. Power (W) $\downarrow$ Aver. Acc. ($\cdots$-2) $\downarrow$ EE pos. error (m) $\downarrow$ EE ori. error (rad) $\downarrow$ Nominal Devi. (m) $\downarrow$
@@ V-C2 Loco-manipulation tracking error
fixed points $0.028\pm 0.019$ $0.089\pm 0.073$ $0.145$
spatial circle $0.048\pm 0.022$ $0.085\pm 0.025$ $0.181$

### v2

@@ IV-A RP: Reward Prioritization
[…]
@@ IV-C Loco-Mani Fusion: Achieving Smooth Loco-Manipulation
$\epsilon_{t}^{\text{ref}}=\max\left\{\epsilon_{0}-vt,0\right\}$ (11)
@@
Action rate $\left|a_{t-1}-a_{t}\right|^{2}$ -0.003
Action smoothness $\left|a_{t-2}-2a_{t-1}+a_{t}\right|^{2}$ -0.001
Collision $1\enspace\text{if collides else}\enspace 0$ -5
Alive $1\enspace\text{if alive else}\enspace 0$ 2.0
Weighted Power $\sum\|\tau\cdot\dot{q}\|\cdot\Omega_{p}$ $3.3\times 10^{-4}$
Weighted Torque $\sum\tau^{2}\cdot\Omega_{\tau}$ $4\times 10^{-5}$
Relative Torques rate $\sum\frac{\|\tau_{t}-\tau_{t-1}|}{\tau_{\text{limit}}}$ 0.1-0.1
@@
$e^{\text{ref}}_{t}$ $=\max\left\{|\epsilon_{t}^{\text{ref}}-\epsilon_{t}|-\gamma,0\right\}$ (13)
@@ V-B Simulation: Ablation Studies
[…]
@@ V-C2 Loco-manipulation tracking error
fixed points $028\pm 0.019 0.0$$089\pm 0.073$ $0.145$
spatial circle $048\pm 0.022 0.0$$085\pm 0.025$ $0.181$